OpenStudy (doc.brown):

Proofread. Can you help me edit my proof. I'm trying to trim the fat off acceleration formulas to explain it better.

OpenStudy (doc.brown):

Acceleration is a change in velocity over time \[a=\frac{\Delta v}{\Delta t}=\frac{v-v_{0}}{t}\] Rearranging for velocity \[v=v_{0}+at\] Rearranging for time \[t=\frac{v-v_{0}}{a}\] For constant acceleration the average velocity is half the difference in velocity, as well as the total distance over time \[v_{av}=\frac{v_{0}+v}{2}=\frac{x-x_{0}}{t}\] Rearranging to show x \[x=x_{0}+(\frac{v_{0}+v)}{2})t\] Substitute v or t \[x=x_{0}+(\frac{v_{0}+\textbf{v})}{2})t\hspace{40pt}x=x_{0}+(\frac{v_{0}+v)}{2})\textbf{t}\] \[x=x_{0}+(\frac{v_{0}+(v_{0}+at)}{2})t\hspace{40pt}x=x_{0}+(\frac{v_{0}+v)}{2})(\frac{v-v_{0}}{a})\] \[x=x_{0}+\frac{2v_{0}t+at^{2}}{2}\hspace{40pt}x=x_{0}+\frac{v_{0}v-v_{0}^{2}+v^{2}-v_{0}v}{2a}\] \[x=x_{0}+v_{0}t+\frac{1}{2}at^{2}\hspace{40pt}v^{2}=v_{0}^{2}+2a(x-x_{0})\]

OpenStudy (doc.brown):

I changed the first line to Acceleration \(a\) is a change \(\Delta\) in velocity \(v\) from the initial velocity \(v_0\) over time \(t\).

OpenStudy (anonymous):

Every equation you used is valid only if acceleration is constant, not just the ones starting with the fourth equation.

OpenStudy (doc.brown):

Thanks, first line is now While acceleration is constant, \(a\) is a change \(\Delta\) in velocity \(v\) from the initial velocity \(v_0\) over time \(t\).

OpenStudy (anonymous):

nice calc-free approach ^_^ This is annoyingly picky, but Delta t only equals t at t_o = 0 :P I'm a stickler. Right under "rearranging to find x" there are a few parenthetical mistypes on the numerator is the only thing I can find to fix. Also, maybe if you make all of the velocity terms (v + v_o) to mirror the form (v - v_o) it'll flow better, and you won't have to do that wonky spread out quadratic 2nd up from the bottom on the left. Last, if you do " \left( and \right) " for parenthesis with fractions \left( \frac{}{} \right) , it looks super slick \[ (\frac{ \textrm{normal parenthesis}}{\textrm{are tiny!!}}) \ \ \ \left( \frac{ \textrm{but these ones are}}{\textrm{super snazzy and profo ~_^}}\right)\] Awesome job all around though!

OpenStudy (doc.brown):

@AllTehMaffs I really am looking for spit and polish, so having someone be annoyingly picky is exactly why I posted. I will implement your changes presently. Thank you so much!

OpenStudy (doc.brown):

I don't want to be stuck with t0 = 0 for a general formula, thanks @simplephysics. But I also don't want to have delta t in all the formulas. How do I shave off the delta?

OpenStudy (anonymous):

I suppose you would have to include a statement at the top specifying that t represents any time interval? \[t = t _{f}-t _{i}\]

OpenStudy (anonymous):

I think t0 has to be 0 for your equation to work and make sense, if not, you wouldn't be talking about a point in time, but about what the speed is at a specific interval delta t equal to t.

OpenStudy (doc.brown):

While acceleration is constant, acceleration \(a\) is a change \(\Delta\) in velocity \(v\), from the initial velocity \(v_0\), over a change in time \(t\) from the initial time \(t_0=0\). \[a=\frac{\Delta v}{\Delta t}=\frac{v-v_{0}}{t-t_0}=\frac{v-v_{0}}{t}\] Rearranging for velocity \[v=v_{0}+at\] Rearranging for time \[t=\frac{v-v_{0}}{a}\] Average velocity is half the difference in velocity, as well as the total distance over time \[v_{av}=\frac{v+v_{0}}{2}=\frac{x-x_{0}}{t}\] Rearranging to show x \[x=x_{0}+\left(\frac{v+v_{0}}{2}\right)t\] Substitute \(\color{red}{v}\) or \(\color{blue}{t}\) \[x=x_{0}+\left(\frac{\color{red}{v}+v_0}{2}\right)t\hspace{75pt}x=x_{0}+\left(\frac{v+v_{0}}{2}\right)\color{blue}{t}\] \[x=x_{0}+\left(\frac{(v_{0}+at)+v_0}{2}\right)t\hspace{40pt}x=x_{0}+\left(\frac{v+v_{0}}{2}\right)\left(\frac{v-v_{0}}{a}\right)\] \[x=x_{0}+\frac{2v_{0}t+at^{2}}{2}\hspace{75pt}x=x_{0}+\frac{v^2+v_{0}v-v_{0}v-v_0^{2}}{2a}\] \[x=x_{0}+v_{0}t+\frac{1}{2}at^{2}\hspace{75pt}v^{2}=v_{0}^{2}+2a(x-x_{0})\]

OpenStudy (doc.brown):

``` While acceleration is constant, acceleration \(a\) is a change \(\Delta\) in velocity \(v\), from the initial velocity \(v_0\), over a change in time \(t\) from the initial time \(t_0=0\). \[a=\frac{\Delta v}{\Delta t}=\frac{v-v_{0}}{t-t_0}=\frac{v-v_{0}}{t}\] Rearranging for velocity \[v=v_{0}+at\] Rearranging for time \[t=\frac{v-v_{0}}{a}\] Average velocity is half the difference in velocity, as well as the total distance over time \[v_{av}=\frac{v+v_{0}}{2}=\frac{x-x_{0}}{t}\] Rearranging to show x \[x=x_{0}+\left(\frac{v+v_{0}}{2}\right)t\] Substitute \(\color{red}{v}\) or \(\color{blue}{t}\) \[x=x_{0}+\left(\frac{\color{red}{v}+v_0}{2}\right)t\hspace{75pt}x=x_{0}+\left(\frac{v+v_{0}}{2}\right)\color{blue}{t}\] \[x=x_{0}+\left(\frac{(v_{0}+at)+v_0}{2}\right)t\hspace{40pt}x=x_{0}+\left(\frac{v+v_{0}}{2}\right)\left(\frac{v-v_{0}}{a}\right)\] \[x=x_{0}+\frac{2v_{0}t+at^{2}}{2}\hspace{75pt}x=x_{0}+\frac{v^2+v_{0}v-v_{0}v-v_0^{2}}{2a}\] \[x=x_{0}+v_{0}t+\frac{1}{2}at^{2}\hspace{75pt}v^{2}=v_{0}^{2}+2a(x-x_{0})\]

OpenStudy (doc.brown):

Should the vectors be bold?

OpenStudy (anonymous):

technically yah. Or you could do /vec \[\vec v \] the bold is fancier though \[ \textbf v\] If you are doing vectors though, you should change your x to something that isn't a cartesian coordinate; s or d something.

OpenStudy (doc.brown):

I've heard people say vee naught. I thought that meant velocity at something-or-other zero \(v_0\). Is it vee oh \(v_o\)?

OpenStudy (anonymous):

It is vee naught, and after posting that I checked all my text books and it's definitely a zero as the subscript, so I'm just crazy. sorry 'bout that ^^

OpenStudy (doc.brown):

I admire a person who researches their own beliefs, and finds them wanting. I wish I could offer you more medals. Thank you, sincerely, for your help.

OpenStudy (anonymous):

I know to derive this in a shorter way.. using vt graph.. if u use vt graph then u can derive them geometrically.. which seems easier cause not much algebra!

OpenStudy (doc.brown):

@Mashy I'd love to read it!

OpenStudy (anonymous):

here it is.. i dunno if i actually made it lengthier :D

OpenStudy (doc.brown):

I stay true to my word, I loved reading that. @Mashy Does the exercise hold true if acceleration isn't constant?

OpenStudy (nincompoop):

same lecture in Gilbert Strang's Physics' book

OpenStudy (doc.brown):

In every physics book. I'm not claiming originality. Just seeking efficiency.

OpenStudy (nincompoop):

not your algebra manipulation, but the derivative

OpenStudy (anonymous):

no.. if acceleration is not a constant, then the graph wouldn't be a straight line!